# All-in-One Transferring Image Compression from Human Perception to Multi-Machine Perception

### v1

@@ Abstract.
Efficiently transferring Learned Image Compression (LIC) model from human perception to machine perception is an emerging challenge in vision-centric representation learning. Existing approaches typically adapt LIC to downstream tasks in a single-task manner, which is inefficient, lacks task interaction, and results in multiple task-specific bitstreams. To address these limitations, we propose an asymmetric adaptor framework that supports multi-task adaptation within a single model. Our method introduces a shared adaptor to learn general semantic features and task-specific adaptors to preserve task-level distinctions. With only lightweight plug-in modules and a frozen base codec, our method achieves strong performance across multiple tasks while maintaining compression efficiency. Experiments on the PASCAL-Context benchmark demonstrate that our method outperforms both Fully Fine-Tuned and other Parameter Efficient Fine-Tuned (PEFT) baselines, and validating the effectiveness of multi-vision transferring.
Image Compression, Machine Learning, PEFT
conference: ; ; ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learning
Figure 1. Our unified model leverages lightweight tuning modules (Gray: frozen base codec; Blue: tuning modules) to efficiently transfer a human-centric codec into a multiple machine vision codec, requiring only a single training and inference process.
[…]
1. Introduction
@@
Image coding for machine (ICM) has become increasingly important in recent years, driven by the fact that more and more visual data is processed by machines rather than viewed by humans. This trend is particularly pronounced with the growing dominance of large-scale vision-language model (VLM) training. The core objective of ICM is to learn compact visual representations that reduce transmission cost while preserving accuracy for downstream machine vision tasks. A straightforward solution is to customize the encoder, decoder, or bitstream in a task-specific manner (Chamain et al., 2021; Bai et al., 2021; Choi and Bajic, 2022; Feng et al., 2022, 2025; Liu et al., 2023a, 2021b; Fischer et al., 2024).
@@
However, such task-specific customization often leads to expensive training, increased bitrate consumption, and substantial parameter overhead. Therefore, motivated by the observation that the most accessible base codecs—optimized for human perception—already produce high-quality reconstructions that preserve most of the semantic content in images, recent works propose to adapt them using parameter-efficient fine-tuning (PEFT) directly. By updating only a small number of parameters, these approaches aim to retain the original human-perception quality of the base codec, while efficiently transferring the latent features and the reconstructed images into a more semantically compact space that is better aligned with downstream machine vision tasks. Despite the efficiency advantages of PEFT-based adaptation, existing methods primarily focus on single-task transfer, treating each task in a completely independent and isolated manner, as shown in Table 1, which incurs several limitations: (1) Each task still requires separately customized training, which is inefficient and hinders scalability as the number of tasks increases; (2) The lack of cross-task interaction prevents the model from learning shared representations across related task domains; (3) Independent adaptation leads to task-specific bitstreams, resulting in additional storage overhead and deployment complexity.
To address the limitations of existing single-task adaptation methods, we propose an asymmetric adaptor framework that enables efficient multi-task transfer within a unified image compression model, as shown in Figure 1. Specifically, to extract task-general and semantically compact features, we introduce a Shared Adaptor, inserted during the Analysis transform (Encoder) and shared across all tasks. This module captures common semantic patterns in a task-agnostic and parameter-efficient manner, serving as a global representation pool. To preserve task-specific information and prevent mutual interference, we introduce a Task Adaptor in the Synthesis transform (Decoder). Each Task Adaptor consists of parallel, task-specific branches that independently modulate features according to the needs of each downstream task. Both components are lightweight, plug-and-play, and trained via a parameter-efficient strategy, ensuring minimal overhead while preserving the base codec’s original compression performance. Together, these modules enable a balance between effective feature sharing and explicit task decoupling, supporting scalable and flexible multi-task adaptation.
@@
  • We propose an asymmetric adaptor framework for efficient multi-task adaptation of pre-trained image codecs, enabling a unified model to support multiple downstream tasks
  • We design two lightweight plug-in modules—Shared Adaptor and Task Adaptor—that jointly enable cross-task feature sharing and task-specific decoupling in a parameter-efficient manner.
  • We demonstrate the effectiveness of our method on the PASCAL-Context benchmark, achieving superior multi-task performance compared to both fully fine-tuned and single-task PEFT baselines.
Table 1. Comparison of representative codec transfer methods in terms of parameter efficiency and multi-task transferability. A method is considered Efficient if it enables adaptation with only a small number of trainable parameters, and supports Multi-Task Transfer if it allows joint optimization across multiple downstream tasks.
Method Efficient Multi-Task Transfer
Full-Finetune
ICMH-Net (Liu et al., 2023a) (MM’23)
TransTIC (Chen et al., 2023) (ICCV’23)
AdaptICMH (Li et al., 2024b) (ECCV’24)
Ours
Figure 2. Overview of our proposed architecture. The backbone follows a typical autoencoder structure from Lu2022-TIC, consisting of an analysis network $g_{a}$ and a synthesis network $g_{s}$. Each encoder and decoder stage includes a convolution layer to adjust feature resolution and a Residual Swin Transformer Block (RSTB) for enhanced representation learning. We introduce two types of adaptor modules: (1) Shared Adaptor (S-Adaptor) is inserted between encoder stages, with dual-branch modulating features in both spatial and frequency domains; (2) Task Adaptor (T-Adaptor) is inserted between decoder stages, composed of $T$ task-specific paths (where $T$ is the number of tasks), each independently processing task-specific features. T-Adaptor outputs are fed into task-specific feature aggregation modules for downstream processing.
[…]
2. Related Works
@@
Learned Image Compression (LIC) was first introduced by Ballé et al. (2017), which adopts an autoencoder-based architecture to perform transform coding in the pixel space. Due to its superior rate-distortion (R-D) performance, LIC has shown great potential as a promising alternative to traditional image compression paradigms. A typical LIC codec consists of three key components: an analysis transform that maps the input image from the high-dimensional pixel space to a compact latent representation; an entropy model that encodes the latent variables into a compressed bitstream; and a synthesis transform that reconstructs the image from the latent space back to the pixel domain. The main research directions in LIC can be broadly categorized into two types. The first focuses on designing more efficient and expressive codec architectures, including more representative analysis and synthesis transforms. These structures have evolved from early CNN-based (Ballé et al., 2017, 2018; Cui et al., 2022; Mentzer et al., 2020; Cheng et al., 2020) designs to Transformer-based (Lu et al., 2021; Li et al., 2024a; Zou et al., 2022; Liu et al., 2023b) models, enabling better modeling capacity. In addition, recent efforts explore user-controllable compression, such as variable-rate coding (Yang et al., 2022; Choi et al., 2019; Li et al., 2024c) and distortion–perception (Agustsson et al., 2023) trade-off control, to enhance flexibility in practical applications. The second line of research focuses on designing more powerful entropy models to better estimate the probability distribution of latent representations. This has evolved from factorized (Ballé et al., 2017) and hyperprior-based (Ballé et al., 2018) models to more advanced autoregressive entropy models (Lee et al., 2019; He et al., 2021; Jiang et al., 2023; Qian et al., 2022; Minnen and Singh, 2020; Minnen et al., 2018). owever, most LIC methods are human-centric, as they are typically optimized using perceptual quality metrics such as MSE or LPIPS. While effective for human viewing, such objectives may not align with the needs of machine vision. In particular, pixel-wise distortion metrics tend to over-allocate bits to visually fine-grained details, while potentially neglecting semantically important structures that are critical for downstream vision tasks
@@
The rapid growth of computer vision applications has led to an increasing portion of visual data being consumed by machines rather than humans—for example, in autonomous driving, traffic monitoring, and visual surveillance. This shift motivates the need for image compression systems optimized for machine vision, aiming to jointly minimize transmission cost while preserving task-relevant information for downstream analysis. A straightforward solution is to define distinct encoder–decoder pairs tailored to each specific task. These approaches can be further categorized into two types: (1) Independent designs (Liu et al., 2022; Chamain et al., 2020; Fischer et al., 2024; Chamain et al., 2021), where each task is handled by a separately trained codec; and (2) Scalable designs (Yan et al., 2021; Feng et al., 2025; Bai et al., 2021), which share part of the model parameters while adapting to different tasks via task-specific components. However, this line of work often leads to multiple task-specific bitstreams and requires dedicated design for each task, resulting in increased system complexity and poor scalability as the number of tasks grows. An alternative line of work aims to extract a unified bitstream for multiple tasks. Specifically, these methods design a task-general encoder to produce semantically compact and generalizable representations that serve as a common source for all tasks. Task-specific outputs are then generated by applying either task-specific decoders (Feng et al., 2022; Liu et al., 2024) or task-guided transformations (Choi and Bajic, 2022) to adapt the shared representation to the needs of each downstream task. However, such codecs designed specifically for multiple vision tasks typically require non-trivial codec architecture design and computationally expensive training, making it difficult to scale to many tasks and challenging to deploy in practice.
@@
3. Method
@@
Given a pre-trained base codec optimized for human perception, our goal is to efficiently transfer it to multiple machine vision tasks within a single unified model and through a single-stage training process. To achieve this, we adopt a multi-task adaptation framework that promotes feature sharing across tasks while preserving task-specific representations through structural decoupling, as illustrated in Figure 2. Furthermore, the training of our framework follows the parameter-efficient fine-tuning (PEFT) paradigm by introducing lightweight plug-in modules, enabling effective adaptation without compromising the original rate-distortion performance of the base codec.
@@
We begin in Section 3.1 by reviewing the base codec architecture and its human perception-based training objective. Then, in Section 3.2, we provide an overview of the key challenges in multi-vision transfer and introduce our asymmetric adaptor-based solution. Next, in Section 3.3, we detail the designs of the adaptation modules we employed, which enable efficient feature sharing and task-specific decoupling. Finally, Section 3.4 describes the joint optimization strategy used to train our model in an end-to-end manner.
3.1. Preliminary
@@
(1) $\mathcal{L}_{rd}=\mathbb{E}_{X\sim p_{X}}[r(\hat{y})+\lambda\cdot\mathcal{D}(X,\hat{X})]$
@@
In this work, we adopt Lu2022-TIC (Lu et al., 2021) as our base codec, which integrates convolutional layers and Transformer blocks in a hybrid architecture. This design enables the model to capture both local and long-range dependencies, supporting effective multi-scale feature extraction. Each encode and decode stage leverages a Residual Swin Transformer Block (RSTB) after a convolutional transformation, formulated as:
(2) $\mathbf{F}_{i}=\text{RSTB}({\text{Conv}(\mathbf{F}_{i-1})})$
@@
3.2. Framework Overview
Given a pre-trained image codec optimized for human perception (e.g., using MSE or LPIPS as the distortion metric), our goal is to adapt it efficiently to multiple downstream vision tasks within a single model. This poses a fundamental challenge: the model must strike a delicate balance between feature sharing and task decoupling. On one hand, different vision tasks often rely on partially overlapping visual cues (e.g., structural boundaries, semantic regions, textures), making shared representations beneficial. Effective feature sharing reduces redundancy, improves generalization, and helps prevent overfitting. On the other hand, excessive coupling may cause negative interference, such as easier tasks dominant more challenging ones. As a result, the architecture must support shared representation learning while preserving sufficient capacity for task-specific adaptation.
To address these challenges, we propose an asymmetric adaptor framework that extends the frozen base codec with two targeted architectural modifications, as illustrated in Figure 2 (Left). First, to enable generalizable feature extraction, we insert a Shared Adaptor (S-Adaptor) into each encode stage and the main path of the decode stage. The S-Adaptor adopts a dual-branch structure that captures both spatial and frequency-domain characteristics, serving as a task-agnostic feature pool for efficient representation sharing. Second, to support task-specific specialization, we insert a Task Adaptor (T-Adaptor) after each decode stage. The T-Adaptor comprises $T$ parallel paths ($T$ is the number of tasks)—one for each task, allowing each task to modulate shared features independently without mutual interference. Finally, the task-specific features from the T-Adaptor form a multi-scale hierarchy, where features at different resolutions encode information at varying levels. These are fused through a lightweight, task-specific aggregation module to generate residuals tailored for each task. he residuals are added to the shared reconstruction $\hat{X}$ to produce the final task-specific output $\hat{X}^{(t)}$ for downstream tasks.
3.3. Adaptation Modules
S-Adaptor Design
To promote feature generalization across tasks, we introduce the Shared Adaptor (S-Adaptor), a lightweight, task-agnostic modulation unit inserted into each encode stage as well as the shared main path of the decode stages. Besides learning shared representations, the S-Adaptor serves as an implicit communication hub, enabling inter-task interaction via a shared representation space without requiring explicit fusion operations.
Motivated by the observation that different vision tasks are sensitive to distinct frequency components (Li et al., 2024b), the S-Adaptor adopts a dual-branch architecture to extract both spatial and frequency-domain features. As illustrated in Figure 2 (Right), given an intermediate feature map $\mathbf{F}_{i}^{\text{in}}\in\mathbb{R}^{H\times W\times C}$ from encode stage $i$, we apply spatial downsampling (omitted in the formulation for simplicity) to reduce computational overhead, followed by two parallel processing streams: (1) The spatial branch employs a depth-wise convolution to capture localized spatial patterns; (2) The frequency branch transforms the feature map into the frequency domain via 2D Fast Fourier Transform (FFT), modulates it with a learnable complex-valued mask $\mathbf{M}_{i}$, and reconstructs it via inverse FFT (IFFT). The outputs from both branches are fused as follows:
@@ 3.3. Adaptation Modules
(3) $\mathbf{F}_{i}^{\text{out}}$ $=\underbrace{\text{DWConv}(\mathbf{F}_{i}^{\text{in}})}_{\text{Spatial Branch}}+\underbrace{\text{IFFT}(\mathbf{M}_{i}\odot\text{FFT}(\mathbf{F}_{i}^{\text{in}}))}_{\text{Frequency Branch}}$
Here, $\text{DWConv}(\cdot)$ denotes depth-wise convolution, and $\odot$ represents element-wise multiplication. $\mathbf{M}_{i}$ is a learnable spectral mask that adaptively emphasizes or suppresses specific frequency components.
T-Adaptor Design
While the S-Adaptor supports shared representation learning, it lacks the ability to tailor features toward task-specific objectives. To address this limitation, we introduce the Task Adaptor (T-Adaptor), which performs task-specific modulation in a decoupled manner. As shown in Figure 2 (middle), the T-Adaptor is inserted after each decode stage and consists of $T$ parallel paths—one per task. Each path follows the same dual-branch architecture as the S-Adaptor but is independently parameterized. Formally, the T-Adaptor at stage $i$ is defined as:
@@
Given a shared feature $\mathbf{G}_{i}\in\mathbb{R}^{H\times W\times C}$ from decode stage $i$, the modulated output for task $t$ is computed as:
(5) $\mathbf{G}_{i}^{(t)}=\text{T-Adaptor}^{(t)}(\mathbf{G}_{i}),\quad t=1,\dots,T$
[…]
Multi-scale Feature Fusion
The task-specific features $\mathbf{G}_{i}^{(t)}$ produced at different decode stages form a multi-scale representation hierarchy. To generate residual information tailored for each task, we aggregate these features using a lightweight multi-scale fusion module. The fused output is then added to the shared reconstruction $\hat{X}$ to obtain the final task-specific prediction:
(6) $\hat{\mathbf{X}}^{(t)}=\hat{X}+\text{Fuse}\left(\left\{\mathbf{G}_{i}^{(t)}\right\}_{i=1}^{3}\right),\quad t=1,\dots,T$
This design ensures that both low-level and high-level task-specific features are effectively integrated into the final reconstruction, enhancing task-adaptive detail refinement. The design of $\text{Fuse}(\cdot)$ is discussed in detail in Section 4.5.
3.4. Optimization Strategy
Our framework is trained end-to-end to jointly optimize compression efficiency and downstream task performance. To quantify the compression efficiency, we adopt the rate-distortion formulation defined in Equation 1, which balances the trade-off between bitrate and reconstruction quality.
To incorporate supervision from downstream tasks, we define the task loss as a weighted combination of task-specific objectives:
(7) $\mathcal{L}_{\text{task}}=\sum_{i=1}^{T}w_{i}\cdot\mathcal{L}_{i},$

### v2

# All-in-One Transfer Image Compression from Human Perception to Multi-Machine Perception

@@ Abstract
Efficiently transferring Learned Image Compression (LIC) model from human perception to machine perception is an emerging challenge in vision-centric representation learning. Existing approaches typically adapt LIC to downstream tasks in a single-task manner, which is inefficient, lacks task interaction, and results in multiple task-specific bitstreams. In this paper, we propose a multi-task adaptation framework that enables transferring a pre-trained base codec to multiple machine vision tasks through a unified model and a single training process. To achieve this, we design an asymmetric adaptation architecture consisting of a task-agnostic encoder adaptation and task-specific decoder adaptation. Furthermore, we introduce two feature propagation modules to facilitate inter-task and inter-scale feature represenation learning. Experiments on PASCAL-Context and NYUD-V2 dataset demonstrate that our method outperforms both Fully Fine-Tuned and other Parameter Efficient Fine-Tuned (PEFT) baselines. Code will be released.
1 Introduction
@@
In recent years, Image Coding for Machine (ICM) has seen a surge in demand, driven by the rapid advancement of deep learning technologies. As a result, an increasing volume of visual data is being processed by machines instead of being viewed by humans. The primary goal of ICM is to learn compact visual representations that minimize transmission cost while preserving high accuracy for downstream vision tasks. To this end, a straightforward solution is to customize the encoder, decoder, or bitstream in a task-specific manner [7, 3, 10, 15, 16, 30, 29, 17].
Method Efficient MT Transfer
Full-Finetune
ICMH-Net [30]
TransTIC [8]
AdaptICMH [24]
Ours
Table 1: Comparison of representative codec transfer methods in terms of parameter efficiency and multi-task transferability. A method is considered Efficient if it enables adaptation with only a small number of trainable parameters, and supports Multi-Task (MT) Transfer if it allows joint optimization across multiple downstream tasks.
@@
However, such task-specific customization often results in costly training procedures, increased bitrate consumption, and significant parameter overhead. To address these limitations, recent studies have turned to a more efficient alternative. Motivated by the observation that widely available base codecs—originally optimized for human perception—can already produce high-quality reconstructions that preserve most of the semantic content in images, these works explore parameter-efficient fine-tuning (PEFT) as a means to adapt such codecs directly for machine vision tasks. By updating only a small number of parameters, PEFT-based methods aim to retain the original human-perception quality of the base codec, while efficiently transferring the latent features and the reconstructed images into a more semantically compact space that is better aligned with downstream machine vision tasks. Despite the efficiency advantages of PEFT-based adaptation, existing methods predominantly focus on single-task transfer, treating each task in a completely independent and isolated manner, as summarized in Table 1, which incurs several limitations: (1) Each task still requires separately customized training, which is inefficient and hinders scalability as the number of tasks grows; (2) The lack of cross-task interaction prevents the model from leveraging shared representations across related task domains; (3) Task-specific adaptations produce separate bitstreams, leading to increased storage requirements and greater complexity in deployment.
To address these limitations, we propose a multi-task adaptation framework that enables transferring a pre-trained base codec to multiple vision tasks using a unified model and a single-stage training process. At the core of our design is a structurally decoupled architecture that introduces parallel, task-specific adaptors to preserve task specialization. Specifically, we redesign each decoding stage of the base codec as a parallel multi-path structure, where each path corresponds to a specific task and leverages an independent adaptor to extract task-specific features. The extracted set of task features is subsequently processed through both inter-task and inter-scale information propagation. First, we introduce a Task Aggregation Module, which integrates features from different tasks and applies a channel selection mechanism to allow each task to access and incorporate complementary information from others. Second, we introduce a Multi-Scale Fusion Module, which applies cross-scale attention to fuse and selectively refine features across different scales. This enables the integration of information from diverse receptive fields.
@@
  • A multi-task adaptation framework that enables transferring a pre-trained base codec to multiple vision tasks within a single model and a single-stage training process.
  • Two modules to enhance representation learning: Task Aggregation Module and Multi-Scale Fusion, for cross-task and cross-scale interaction, respectively.
  • We validate our approach on two widely-used multi-task learning benchmarks, PASCAL-Context and NYUD-V2, demonstrating superior performance.
2 Related Works
@@
Learned Image Compression (LIC) was first introduced by Ballé et al. [4], which adopts an autoencoder-based architecture to perform transform coding in the pixel space. Due to its superior rate-distortion (R-D) performance, LIC has shown great potential as a promising alternative to traditional image compression paradigms. A typical LIC codec consists of three key components: an analysis transform that maps the input image from the high-dimensional pixel space to a compact latent representation; an entropy model that encodes the latent variables into a compressed bitstream; and a synthesis transform that reconstructs the image from the latent space back to the pixel domain. The main research directions in LIC can be broadly categorized into two types. The first focuses on designing more efficient and expressive codec architectures, including more representative analysis and synthesis transforms. These structures have evolved from early CNN-based [4, 5, 12, 34, 9] designs to Transformer-based [32, 23, 43, 27] models, enabling better modeling capacity. In addition, recent efforts explore user-controllable compression, such as variable-rate coding [42, 11, 22] and distortion–perception [2] trade-off control, to enhance flexibility in practical applications. The second line of research focuses on designing more powerful entropy models to better estimate the probability distribution of latent representations. This has evolved from factorized [4] and hyperprior-based [5] models to more advanced autoregressive entropy models [21, 18, 20, 38, 35, 36]. owever, most LIC methods are human-centric, as they are typically optimized using perceptual quality metrics such as MSE or LPIPS. While effective for human viewing, such objectives may not align with the needs of machine vision. In particular, pixel-wise distortion metrics tend to over-allocate bits to visually fine-grained details, while potentially neglecting semantically important structures that are critical for downstream vision tasks
@@
The rapid growth of computer vision applications has led to an increasing portion of visual data being consumed by machines rather than humans—for example, in autonomous driving, traffic monitoring, and visual surveillance. This shift motivates the need for image compression systems optimized for machine vision, aiming to jointly minimize transmission cost while preserving task-relevant information for downstream analysis. A straightforward solution is to define distinct encoder–decoder pairs tailored to each specific task. These approaches can be further categorized into two types: (1) Independent designs [26, 6, 17, 7], where each task is handled by a separately trained codec; and (2) Scalable designs [41, 16, 3], which share part of the model parameters while adapting to different tasks via task-specific components. However, this line of work often leads to multiple task-specific bitstreams and requires dedicated design for each task, resulting in increased system complexity and poor scalability as the number of tasks grows. An alternative line of work aims to extract a unified bitstream for multiple tasks. Specifically, these methods design a task-general encoder to produce semantically compact and generalizable representations that serve as a common source for all tasks. Task-specific outputs are then generated by applying either task-specific decoders [15, 28] or task-guided transformations [10] to adapt the shared representation to the needs of each downstream task. However, such codecs designed specifically for multiple vision tasks typically require non-trivial codec architecture design and computationally expensive training, making it difficult to scale to many tasks and challenging to deploy in practice.
@@
Figure 1: Overview of our proposed architecture. The backbone adopts a standard autoencoder design based on Lu2022-TIC, consisting of an encoder and a decoder. The encoder follows a single-path structure, with a task-agnostic Feature Extraction Adaptor inserted after each encoding stage. In contrast, the decoder follows a multi-path design, where each path contains a task-specific Feature Extraction Adaptor. The resulting task features are further processed by the Task Aggregation Module and the Multi-Scale Fusion Module to enable both inter-task and inter-scale information propagation.
3 Method
@@
Given a pre-trained base codec optimized for human perception, our goal is to efficiently transfer it to multiple machine vision tasks within a single unified model and through a single-stage training process. This poses a fundamental challenge: the model must strike a delicate balance between task feature sharing and decoupling. On the one hand, it should leverage the overlapping visual cues across vision tasks; on the other hand, it must avoid excessive coupling, which can lead to negative task interference. To achieve this, we adopt a multi-task adaptation framework that promotes feature sharing across tasks while preserving task-specific representations through structural decoupling. Furthermore, the training of our framework follows the parameter-efficient fine-tuning (PEFT) paradigm by introducing lightweight plug-in modules, enabling effective adaptation without compromising the original rate-distortion performance of the base codec.
@@
We begin by reviewing the architecture of the base codec and its training objective, which is optimized for human perception. We then present an overview of our proposed multi-task adaptation framework. Subsequently, we detail the design of the adaptation components, including the Task Aggregation Module and the Multi-Scale Fusion Module, which facilitate inter-task and inter-scale feature interaction. Finally, we describe the joint optimization strategy used to train our model.
3.1 Preliminary
@@
$\mathcal{L}_{rd}=\mathbb{E}_{x\sim p_{x}}[r(\hat{y})+\lambda\cdot\mathcal{D}(x,\hat{x})]$ (1)
@@
In this work, we adopt Lu2022-TIC [32] as our base codec, which integrates convolutional layers and Transformer blocks in a hybrid architecture. This design enables the model to capture both local and long-range dependencies, supporting effective multi-scale feature extraction. Each encode and decode stage leverages a Residual Swin Transformer Block (RSTB) after a convolutional transformation, formulated as:
$\mathbf{F}_{i}=\text{RSTB}({\text{Conv}(\mathbf{F}_{i-1})})$ (2)
@@
3.2 Framework Overview
As shown in Figure 1, given a pre-trained image codec optimized for human perception (e.g., using MSE or LPIPS as the distortion metric), our goal is to adapt it efficiently to multiple downstream vision tasks within a single model.
To achieve this, we propose an asymmetric adaptation framework that extends the frozen base codec with two targeted architectural modifications. In the encoding stage, our goal is to enable generalizable feature extraction. To this end, we insert a single lightweight and task-agnostic Feature Extraction adaptor into each encoding stage (we deliberately avoid task-specific modulation to prevent multiple bitstreams). In the decoding stage, to support task-specific specialization, we insert multiple parallel Feature Extraction adaptors to independently extract task-specific features for each task. These features are then passed to the Task Aggregation Module, where information is exchanged across tasks to formulate a shared representation and enhance each task’s feature by incorporating complementary cues from other tasks. Finally, the shared feature is propagated as the primary representation to the subsequent decoding stage. Task-specific features produced at different stages are organized into a multi-scale hierarchy, which is further processed by a Multi-Scale Fusion Module. The fused task representations are then forwarded to task-specific decode heads to generate the final predictions for each task.
3.3 Task Aggregation Module
Feature Extraction Adaptor
Feature Extraction (FE) adaptor is inserted into each encoding stage and is also placed in parallel along each task-specific path in the decoding stage. Within our framework, it serves as a basic unit for extracting informative representations. In this paper, we directly adopt the FE adaptor design proposed [24] as our FE adaptor, which employs a dual-branch architecture to simultaneously capture spatial and frequency-domain features. Given an intermediate feature map $\mathbf{F}^{\text{in}}\in\mathbb{R}^{H\times W\times C}$, the adaptor processes it through two parallel streams:
@@ 3.3 Task Aggregation Module
$\mathbf{F}^{\text{out}}$ $=\underbrace{\text{DWConv}(\mathbf{F}^{\text{in}})}_{\text{Spatial Branch}}+\underbrace{\text{IFFT}(\mathbf{M}\odot\text{FFT}(\mathbf{F}^{\text{in}}))}_{\text{Frequency Branch}}$ (3)
@@
Task Aggregation Module (TAM)
Each decoding stage consists of $N$ parallel Feature Extraction (FE) adaptors, each responsible for generating task-specific features for the corresponding task. To enable cross-task information propagation, these features are passed into a Task Aggregation Module (TAM), where inter-task relationships and dependencies are captured. Within the TAM, task-specific features are further refined by selectively incorporating relevant information from other tasks. In particular, the TAM recieve as input $N$ task-specific features, each of shape $h\times w\times C$, denoted as $\{F_{k}\}_{k=1}^{N}$. These features are first concatenated along the channel dimension to form a unified representationand $\mathbf{F}_{\text{cat}}\in\mathbb{R}^{h\times w\times(NC)}$. $\mathbf{F}_{\text{cat}}$ is then passed through a transformation module $\Phi(\cdot)$, which consists of linear and nonlinear layers followed by a softmax activation, producing a channel-wise attention mask of the same shape. The attention mask is used to reweight the concatenated features via element-wise multiplication, resulting in a shared feature representation:
$\mathbf{F}_{\text{cat}}$ $=\text{Concat}(\mathbf{F}_{1},\ldots,\mathbf{F}_{N})$ (4)
$\mathbf{F}_{\text{shared}}$ $=\mathbf{F}_{\text{cat}}\odot\text{Softmax}\left(\Phi(\mathbf{F}_{\text{cat}})\right)$ (5)
The resulting shared feature follows two propagation paths. First, it is propagated through the main path of the subsequent decoding stages. Second, it simultaneously serves as residual information to inject cross-task knowledge into each task-specific representation. Specifically, the shared feature is directly added to each task-specific feature to extract complementary information from other tasks and enhance the original representations. The refined features are then passed to the multi-scale fusion module, which will be described in the following section.
3.4 Multi-Scale Fusion
While the TAM within each decoding stage enables task interactions, it remains limited to a certain scale. However, prior work [39] has shown that task affinity patterns are scale-dependent, which means the mutual influence between relevant tasks is not necessarily preserved across different scales. Thus, we propose a Multi-Scale Fusion (MSF) module that explicitly incorporates cross-scale information propagation. As illustrated in Figure 1 (right), for individual task, the MSF module takes as input a set of multi-scale features $\{\mathbf{F}_{s}\}_{s=1}^{3}$, obtained from different decoding stages. We treat features with large and medium receptive fields, $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$ as the query and key, and the feature with a the smallest receptive field, $\mathbf{F}_{3}$, as value, to formulate the query-key-value pair. To achieve this, we first align the spatial resolutions of $\mathbf{F}_{1}$, $\mathbf{F}_{2}$, $\mathbf{F}_{3}$, using bilinear interpolation. Then, we apply three separate convolutional layers to extract the query and key from the concatenated features of $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$, and the value from $\mathbf{F}_{3}$. Finally, a cross-scale attention module is employed to produce representations that integrate information across different receptive field scales. Formally, this process is defined as:
$\mathbf{Q},\mathbf{K}=\text{Conv}_{q,k}([\mathbf{F}_{1};\mathbf{F}_{2}]),\quad\mathbf{V}=\text{Conv}_{v}(\mathbf{F}_{3})$ (6)
$\mathbf{F}_{\text{out}}=\text{Softmax}\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{\top}}{\sqrt{d}}\right)\cdot\mathbf{V}$ (7)
Here, $\text{Conv}_{q,k}(\cdot)$ and $\text{Conv}_{v}(\cdot)$ are $\times 1$ 1 convolutions for query/key and value projection, [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes channel-wise concatenation, and $d$ is the channel dimension used for scaling.
3.5 Optimization Strategy
Our framework is trained end-to-end to jointly optimize compression efficiency and downstream task performance. To quantify the compression efficiency, we adopt the rate-distortion formulation defined in Equation 1, which balances the trade-off between bitrate and reconstruction quality. To incorporate supervision from downstream tasks, we define the task loss as a weighted combination of task-specific objectives:
$\mathcal{L}_{\text{task}}=\sum_{i=1}^{T}w_{i}\cdot\mathcal{L}_{i},$ (8)